In 2019, X Company's profit function was 0.30R - $91,400, where R is revenue. In 2020, the relationship between revenue and variable costs will not change, but fixed costs will increase by $20,108. Assuming a tax rate of 40%, what will revenue have to be in order for X Company to earn $44,700 after taxes in 2020?

Tries 0/3
Answer #1

Tax rate = 40%

Profit after tax = $44,700

Profit before tax = Profit after tax/ (1- tax rate)

= 44,700/ (1-0.40)

= 44,700/0.6

= $74,500

Increase in fixed cost = $20,108

In the profit function, 0.30R - $91,400, $91,400 is the fixed cost.

fixed cost will become = 91,400+20,108

= $111,508

Profit function = 0.30R - 111,508

74,500 = 0.30R - 111,508

R = 620,027

Hence, revenue of $620,027 is needed to earn profit of $44,700, after taxes.

Kindly comment if you need further assistance. Thanks‼!